### v1

@@ Results.
We employ PSNR, SSIM [50] and LPIPS [56] as metrics for novel view synthesis. We compare our method against the state-of-the-art large-scale NeRF-based methods: Mega-NeRF [46], Switch-NeRF [31] [^3] https://github.com/MiZhenxing/Switch-NeRF, and 3DGS-based methods: 3DGS [18] [^4] https://github.com/graphdeco-inria/gaussian-splatting, VastGaussian [22], Hierarchy-GS [19]. For Mega-NeRF and Switch-NeRF, we use the officially provided checkpoints on 8 blocks for evaluation. The results of Hierarchy-GS are cited from the original paper since its code has not been released during this work.
We present the quantitative visual quality results in Table 1. The training time and rendering efficiency are also provided in Table 2. As shown in the tables, methods based on 3D Gaussians achieved better results than NeRF-based methods. Although NeRF-based methods are comparable to 3DGS methods in PSNR, the rendered images lack details in such large-scale scenes. This is also validated from Fig. 4 and Fig. 5. Moreover, NeRF-based methods are much slower than 3D-GS-based methods and take longer time to train – even if they are trained distributedly. Notably, our method achieves the best results in almost all the scenes. The original 3DGS has comparable results to us. However, it takes 68×6$\sim$ 8$\times$6 ∼ 8 × more training time than our method. We also emphasize that we build a strong baseline of 3DGS for fair comparison: the densification interval is 200 iterations, which is 8 times more frequent than the 3DGS baseline in VastGaussian; the training iteration is 500K (in comparison, the training iteration of the 3DGS baseline is 450K in VastGaussian). VastGaussian trains 3DGS faster than our method. This is because our method requires additional time for data transmission. However, our method achieves better rendering quality than VastGaussian. Moreover, the data transmission time does not become the bottleneck of our method due to our balanced scene splitting method. Particularly, each block has a similar training time and the master node does not need to wait for a long time for the fat nodes to finish its job.
@@
We ablate the effectiveness of our method and present the results in Table 4. Our method without applying the 3D Gaussian consensus is denoted as w.o. CS, our method without adopting the self-adaptation of penalty parameters is denoted as w.o. SD, our method without adopting the over-relaxation is denoted as w.o. OR. As shown in the table, the performance drastically drops without the ADMM consensus step. Furthermore, the results without applying the self-adaptation of penalty parameters is about 1.5 dB lower than the full model in PSNR. The model without applying over-relaxation is comparable to the full model in SSIM and LPIPS but has lower PSNR. We thus employ over-relaxation in our method. We also present the qualitative differences in Fig. 6, and it clearly shows the full model has better quality in the rendered images and geometries.
@@ Appendix A Appendix
We denote the average of a vector with an overline. Then the consensus step for Eq. (8b) can be rewritten by
@@
Moreover, averaging both sides of Eq. (8c) for the dual variables gives
@@ A.7 More Qualitative Results
Refer to captionRefer to caption
Figure 7: More qualitative results on the mill19 dataset and the UrbanScene3D dataset. Zoom in for best view.
@@ A.8 Limitations and Future Work
Our method can distributedly train 3D GS on large-scale scenes. However, it brings additional communication overheads to the system. Fortunately, we found the communication overhead did not slow down the training performance. This is due to our balanced splitting algorithm minimizing the synchronization time when there is a need to consensus the local 3D Gaussians from all blocks. Moreover, we pruned unnecessary small 3D Gaussians to reduce further the number of 3D GS, which also reduced the communication overhead.
Though our method can train large-scale scenes efficiently, the GPU memory requirement can still be an issue. This is because when we zoom out to capture larger areas, more 3D Gaussians are included in the rasterization step, which consumes more GPU memory. Our future work will consider introducing the level-of-details (LOD) technique into our distributed training pipeline. Like existing LOD GS methods [37, 40], the LOD can be utilized to reduce the number of 3D Gaussians that are far away from the cameras.
@@ A.9 More Discussions
(1) VastGaussian [22] focuses mostly on the data splitting strategy and guaranteeing the consistency of different 3D Gaussians can also be a challenge in different scenes. Our method, however, ensured the consistency of the shared 3D Gaussians through the 3D Gaussian consensus with only a quite simple splitting approach. Nonetheless, the data splitting approach in VastGaussian can be introduced into our framework to enhance the robustness. Moreover, the decoupled appearance encoding can also be applied to our method to reduce floaters. Therefore, our work and VastGaussian are complementary to each other. (2) Hierarchy-GS [19] also trains 3DGS distributedly in different grid cells. However, Hierarchy-GS focuses more on the rendering speed by designing a hierarchical tree structure for 3D Gaussians. During training, the generated hierarchy is loaded and optimized by randomly selecting a training view and a target granularity. After training all individual chunks, Hierarchy-GS first runs a coarse initial optimization on the full dataset and adds an auxiliary skybox. This coarse model serves as a minimal basis for providing backdrop details for parts of the scene that are outside of a chunk. When consolidating different chunks, a 3D Gaussian primitive is deleted if it is associated to-but outside-chunk and is closer to another chunk to maintain the consistency of 3D Gaussians. The naive pruning approach neglects the gradient flow of the 3D Gaussians to other 3D Gaussians that are inside only in an associated chunk. Therefore, its rendering quality can be worse than VastGaussian and our method. Nonetheless, its hierarchical LOD approach can still serve as a good complement to our method and VastGaussian.
@@
DoGaussian, VastGaussian, and Hierarchy-GS focus on different parts of training 3D GS on large-scale scenes and therefore are complementary to each other. These methods can be consolidated into a more robust and scalable system to open a new world in city-scale 3D reconstruction. Our method, however, requires an additional master node for controlling the training the 3DGS, which consumes more GPU resources.

### v2

@@ Results.
We employ PSNR, SSIM [55] and LPIPS [61] as metrics for novel view synthesis. We compare our method against the state-of-the-art large-scale NeRF-based methods: Mega-NeRF [51], Switch-NeRF [35] [^3] https://github.com/MiZhenxing/Switch-NeRF, and 3DGS-based methods: 3DGS [21] [^4] https://github.com/graphdeco-inria/gaussian-splatting, Fed3DGS [48], VastGaussian [26], Hierarchy-GS [22]. For Mega-NeRF and Switch-NeRF, we use the officially provided checkpoints on 8 blocks for evaluation. The results of Hierarchy-GS are cited from the original paper since its code has not been released during this work.
We present the quantitative visual quality results in Table 1. The training time and rendering efficiency are also provided in Table 2. As shown in the tables, methods based on 3D Gaussians achieved better results than NeRF-based methods. Although NeRF-based methods are comparable to 3DGS methods in PSNR, the rendered images lack details in such large-scale scenes. This is also validated from Fig. 4 and Fig. 5. Moreover, NeRF-based methods are much slower than 3DGS-based methods and take longer time to train – even if they are trained distributedly. Notably, our method achieves the best results in almost all the scenes. The original 3DGS has results comparable to ours. However, it takes 68×6$\sim$ 8$\times$6 ∼ 8 × more training time than our method. We also emphasize that we build a strong baseline of 3DGS for fair comparison: the densification interval is 200 iterations, which is 8 times more frequent than the 3DGS baseline in VastGaussian; the training iteration is 500K (in comparison, the training iteration of the 3DGS baseline is 450K in VastGaussian). VastGaussian trains 3DGS faster than our method. This is because our method requires additional time for data transmission. However, our method achieves better rendering quality than VastGaussian. Moreover, the data transmission time does not become the bottleneck of our method due to our balanced scene splitting method. Particularly, each block has a similar training time and the master node does not need to wait for a long time for the fat nodes to finish its job.
To further show the applicability of our method to larger-scale scenes, we evaluated our method on the $2.7\text{km}^{2}$ Small City scene in the MatrixCity [24] dataset, which contains $5,620$ training views and $741$ validation views. We early terminated the training of the original 3DGS since it did not finish the training within two days. VastGaussian failed on this dataset since two blocks produce no 3D Gaussian primitives due to the imbalanced splitting. From Table 3, our method achieved the best results in rendering quality. The visual qualitative results are shown in Fig. 6.
@@
Table 3: Quantitative results of novel view synthesis on the MatrixCity [24] dataset. $\uparrow$: higher is better, $\downarrow$: lower is better. The red, orange and yellow colors respectively denote the best, the second best, and the third best results.
Scenes aerial street
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time $\downarrow$ Points Mem $\downarrow$ FPS $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time $\downarrow$ Points Mem $\downarrow$ FPS $\uparrow$
3DGS 27.36 0.818 0.237 47:40 11.9 6.31 45.57 20.03 0.643 0.650 14:24 1.85 2.33 193.32
$\text{VastGaussian}^{\dagger}$ 28.33 0.835 0.220 05:53 12.5 6.99 40.04 - - - - - - -
DOGS 28.58 0.847 0.219 06:34 10.3 5.82 48.34 21.61 0.652 0.649 02:33 2.37 2.89 180.51
@@
We ablate the effectiveness of our method and present the results in Table 5. Our method without applying the 3D Gaussian consensus is denoted as w.o. CS, our method without adopting the self-adaptation of penalty parameters is denoted as w.o. SD, our method without adopting the over-relaxation is denoted as w.o. OR. As shown in the table, the performance drastically drops without the ADMM consensus step. Furthermore, the results without applying the self-adaptation of penalty parameters is about 1.5 dB lower than the full model in PSNR. The model without applying over-relaxation is comparable to the full model in SSIM and LPIPS but has lower PSNR. We thus employ over-relaxation in our method. We also present the qualitative differences in Fig. 7, and it clearly shows the full model has better quality in the rendered images and geometries. We include more qualitative results in Fig. 8 to show the importance of the consensus step. We can observe that the distributed training presents noisy results without the consensus step. From the two bottom-right figures, we can observe obvious artifacts along the block boundary without the consensus step.
@@ 5 Conclusion
Acknowledgement. This research / project is supported by the National Research Foundation (NRF) Singapore, under its NRF-Investigatorship Programme (Award ID. NRF-NRFI09-0008). Yu Chen is also partially supported by a Google PhD Fellowship.
@@ Appendix A Appendix
We denote the average of a vector with an overline. Then the consensus step for Eq. (8b) can be rewritten by:
@@
Moreover, averaging both sides of Eq. (8c) for the dual variables gives:
@@ A.7 More Qualitative Results
We present more qualitative results of our method in Fig. 9 and Fif. 10. We provide qualitative comparisons with VastGaussian in areas where blocks overlap in Fig. 10. Both methods produce fairly consistent results. However, our method presents higher fidelity rendering results than VastGaussian near the splitting boundary, which also validated the effectiveness and importance of the consensus step.
Figure 9: More qualitative results on the mill19 dataset and the UrbanScene3D dataset. Zoom in for the best view.
Figure 10: More qualitative comparisons of our method and state-of-the-art methods.
Figure 11: Qualitative results near scene boundary on the UrbanScene3D dataset. Top-Right: VastGaussian. Bottom-Right: our method.
@@ A.8 Limitations and Future Work
Our method can distributedly train 3D GS on large-scale scenes. However, it brings additional communication overheads to the system. Fortunately, we found the communication overhead did not slow down the training performance. This is due to our balanced splitting algorithm minimizing the synchronization time when there is a need to consensus the local 3D Gaussians from all blocks. Moreover, we pruned unnecessary small 3D Gaussians to further reduce the number of 3D GS, which also reduced the communication overhead.
Though our method can train large-scale scenes efficiently, the GPU memory requirement can still be an issue. This is because when we zoom out to capture larger areas, more 3D Gaussians are included in the rasterization step, which consumes more GPU memory. Our future work will consider introducing the level-of-details (LOD) technique into our distributed training pipeline. Similar to existing LOD GS methods [41, 44], LOD can be utilized to reduce the number of 3D Gaussians that are far away from the cameras.
@@ A.9 More Discussions
Training and Waiting Time of Each Block.
We tested the time cost from transferring the data to the master node to receiving data from the master node for each slave node on the Campus dataset. The mean and variance of time are respectively 5.63 seconds and 0.75 seconds each time. The low variance indicates that our method can balance the training time well. We argue that the data transfer time of our method can be kept constant since we can always control the number of local Gaussians to a constant number (e.g. $<=6000,000$ 3D Gaussian primitives) with enough GPUs, no matter how large the scene is, since the data transfer between different slave nodes and the master node is executed distributedly instead of sequentially.
@@
1) VastGaussian [26] focuses mostly on the data splitting strategy and guaranteeing the consistency of different 3D Gaussians can also be a challenge in different scenes. However, our method ensured the consistency of the shared 3D Gaussians through the 3D Gaussian consensus with only a quite simple splitting approach. Nonetheless, the data splitting approach in VastGaussian can be introduced into our framework to enhance the robustness. Moreover, the decoupled appearance encoding can also be applied to our method to reduce floaters. Our work and VastGaussian are thus complementary to each other. 2) Hierarchy-GS [22] also trains 3DGS distributedly in different grid cells. However, Hierarchy-GS focuses more on the rendering speed by designing a hierarchical tree structure for 3D Gaussians. During training, the generated hierarchy is loaded and optimized by randomly selecting a training view and a target granularity. After training all individual chunks, Hierarchy-GS first runs a coarse initial optimization on the full dataset and adds an auxiliary skybox. This coarse model serves as a minimal basis for providing backdrop details for parts of the scene that are outside of a chunk. When consolidating different chunks, a 3D Gaussian primitive is deleted if it is associated to-but outside-chunk and is closer to another chunk to maintain the consistency of 3D Gaussians. The naive pruning approach neglects the gradient flow of the 3D Gaussians to other 3D Gaussians that are inside only in an associated chunk. Therefore, its rendering quality can be worse than VastGaussian and our method. Nonetheless, its hierarchical LOD approach can still serve as a good complement to our method and VastGaussian.
@@
DOGS, VastGaussian and Hierarchy-GS focus on different parts of training 3D GS on large-scale scenes and therefore are complementary to each other. These methods can be consolidated into a more robust and scalable system to open a new world in city-scale 3D reconstruction. Our method, however, requires an additional master node for controlling the training of the 3DGS, which consumes more GPU resources.